OpenStudy (anonymous):

Solve: 6(x − 2) + 3(x − 2) = 18

OpenStudy (alexwee123):

distribute the 6 and 3 first

mathslover (mathslover):

\[\large{6(x-2)+3(x-2)=18}\] try to take (x-2) common

OpenStudy (anonymous):

6x - 12 + 3x -12 = 18

mathslover (mathslover):

no no try it again : 3(x-2)=3(x)-3(2)

OpenStudy (anonymous):

ya.. go on

OpenStudy (anonymous):

3*2=6

OpenStudy (alexwee123):

3(-2)=-6 not -12 :/

mathslover (mathslover):

yes .

OpenStudy (anonymous):

huh

OpenStudy (alexwee123):